# Neural network learns low-dimensional polynomials with SGD near the information-theoretic limit

### v1

@@ 2.3 Challenges in Establishing Statistical Guarantees
Importantly, the analysis in [DTA+24] does not lead to concrete learnability guarantees for the class of single-index polynomials for the following reasons: $(i)$ it is not clear if an appropriate nonlinear transformation that lowers the information exponent can always be extracted from SGD with reused batch, and $(ii)$ the weak recovery guarantee may not translate to a sample complexity for the trained NN to achieve small generalization error. We elaborate these technical challenges below.
@@
To show weak recovery, [DTA+24, Definition 3.1] assumed that the student activation $\sigma$ can reduce the information exponent of the labels to $1$; while a few examples are given (e.g., odd-degree Hermite polynomials), the existence of such transformations in SGD is not guaranteed:
  • The label transformation employed in prior SQ algorithms [CM20] is based on thresholding, which reduces the information exponent to $2$ for any polynomial $\sigma_{*}$; however, isolating such function from SGD updates on the squared loss is challenging. Instead, we show in Proposition 5 that monomial transformation suffices, and it can be extracted from SGD via Taylor expansion.
  • If the link function $\sigma_{*}$ is even, then its information exponent after arbitrary nonlinear transformation is at least $2$; such functions are predicted not be not learnable by SGD in the $n\asymp d$ regime [DTA+24]. To handle this setting, we analyze the SGD update up to $\mathrm{polylog}(d)$ time, at which a nontrivial overlap can be established by a Grönwall-type argument similar to [BAGJ21]. This recovers results on phase retrieval when $\sigma_{*}(z)=z^{2}$ which requires $n=\Omega(d\log d)$.
@@
Note that weak recovery (i.e., $|\langle\boldsymbol{w},\boldsymbol{\theta}\rangle|>\varepsilon$ for some small constant $\varepsilon>0$) is generally insufficient to establish low generalization error of the trained NN. Therefore, we need to show that starting from a nontrivial overlap, subsequent gradient steps can achieve strong recovery of the index features (i.e., $|\langle\boldsymbol{w},\boldsymbol{\theta}\rangle|>1-\varepsilon$), despite the link misspecification. After the first-layer parameters align with the target function, we can train the second-layer parameters with SGD to learn the link function $\sigma_{*}$ with the aid of random bias units.
3 SGD Achieves Almost-linear Sample Complexity
We consider the learning of single-index polynomials with degree $q$ and information exponent $p$:
@@ 3.1 Training Algorithm
We train the following two-layer neural network with $N$ neurons using SGD to minimize the squared loss:
@@
  • The most important part in Phase I of Algorithm 1 is the reuse of the same minibatch in the gradient computation. Specifically, we sample a fresh batch of training examples in every two GD steps; this enables us to extract non-CSQ terms from two consecutive gradient updates outlined in (2.2).
  • We introduce an interpolation step between the current and previous iterates with hyperparameter $\xi$ to stabilize the training dynamics; this resembles a negative momentum term often seen in optimization algorithms [AZ18, ZLBH19]; the role of this interpolation step is discussed in Section 4.2. We also use a projected gradient update ~$\tilde{\nabla}_{\boldsymbol{w}}\mathcal{L}(\boldsymbol{w})=(\boldsymbol{I}_{d}-\boldsymbol{w}^{2t}{\boldsymbol{w}^{2t}}^{\top})\nabla_{\boldsymbol{w}}\mathcal{L}(\boldsymbol{w})$ for steps $2t$ and $12t+$ 1, where $\nabla_{\boldsymbol{w}}$ is the Euclidean gradient. Similar use of projection also appeared in [DNGL23, AAM23].
@@
Input : Learning rates $\eta^{t}$, momentum parameters $\{\xi_{j}^{t}\}_{j=1}^{N}$, number of steps $T_{1},T_{2}$,
$\ell_{2}$ regularization $\lambda$.
@@
3Initialize $\boldsymbol{w}^{0}_{j}\sim\mathbb{S}^{d-1}(1)$, $a_{j}\sim\mathrm{Unif}\{\pm r_{a}\}$.
@@
6             if $t>0$ is even then
7                   Draw i.i.d. sample $(\boldsymbol{x},y)$.
8                   Interpolate ${\boldsymbol{w}}^{t}_{j}\leftarrow\boldsymbol{w}^{t}_{j}-\xi_{j}^{t}(\boldsymbol{w}^{t}_{j}-\boldsymbol{w}^{t-2}_{j})$.
9                   Normalize $\boldsymbol{w}^{t}_{j}\leftarrow\boldsymbol{w}^{t}_{j}/\|\boldsymbol{w}^{t}_{j}\|$.
10            $\boldsymbol{w}^{t+1}_{j}\leftarrow\boldsymbol{w}^{t}_{j}-\eta^{t}\tilde{\nabla}_{\boldsymbol{w}}(f_{\boldsymbol{\Theta}}(\boldsymbol{x})-y)^{2},\quad(j=1,\dots,N)$.
11
12
13Initialize $b_{j}\sim\mathrm{Unif}([-C_{b},C_{b}])$.
14Phase II: SGD on second-layer parameters
15       $\hat{\boldsymbol{a}}\leftarrow\mathrm{argmin}_{\boldsymbol{a}\in\mathbb{R}^{N}}\frac{1}{T_{2}}\sum_{i=1}^{T_{2}}\mathopen{}\mathclose{{}\left(f_{\boldsymbol{\Theta}}(\boldsymbol{x}_{i})-y_{i}}\right)^{2}+\lambda\|\boldsymbol{a}\|^{2}.$
16
@@
3.2 Convergence and Statistical Complexity
@@
We require the activation function to be a polynomial $\sigma(z)=\sum_{i=0}^{C_{\sigma}}\beta_{i}\mathsf{He}_{i}(z)$ and its degree $C_{\sigma}$ to be sufficiently large so that $C_{\sigma}\geq C_{q}$ holds ($C_{q}$ is defined in Proposition 5). For all $2\leq\ell\leq C_{\sigma}$ and $k=0,1$, we assume that $H\big{(}\sigma^{(\ell)}(\sigma^{(1)})^{\ell-1};k\big{)}\neq 0$.
More precisely, a given $\sigma_{*}$, the above condition only needs to be met for one pair of $(k,\ell)$, as specified in Appendix B.1. We show that this condition is satisfied for a wide range of polynomials with degree $C_{\sigma}$.
@@
Given any $\ell\geq 2$ and $k\geq 0$. For $C_{\sigma}\geq\frac{2\ell+k-1}{\ell}$, if we choose $\{\beta_{i}\}_{i=0}^{C_{\sigma}}$ where each $\beta_{i}$ is randomly drawn from some interval $[a_{i},b_{i}]$ with $a_{i}<b_{i}$, then $H(\sigma^{(\ell)}(\sigma^{(1)})^{\ell-1};k)\neq 0$ with probability one.
The following theorem states that $n=\tilde{\Theta}(d)$ samples are sufficient to achieve weak recovery.
@@
Under Assumptions 1 and 2, for suitable choices of hyperparameters $\eta^{t}=\tilde{O}_{d}(Nd^{-1})$ and $\xi_{j}^{t}=O(1)$, there exists constant $C(q)$ such that after Phase I of Algorithm 1 is run for $2T_{1,1}=C(q)\cdot d\mathrm{polylog}(d)$ steps, with high probability, there exists a subset of neurons $\boldsymbol{w}_{j}^{2T_{1}}\in\mathcal{W}$ with $|\mathcal{W}|=\tilde{\Theta}(N)$ such that $\big{|}\langle\boldsymbol{w}_{j}^{2T_{1}},\boldsymbol{\theta}\rangle\big{|}>c$ for some $c\gtrsim 1/\mathrm{polylog}(d)$.
@@
Recall the Hermite expansions $\sigma_{*}(z)=\sum_{i=p}^{q}\alpha_{i}\mathsf{He}_{i}(z)$, $\sigma_{j}(z)=\sum_{i=0}^{C_{\sigma}}\beta_{j,i}\mathsf{He}_{i}(z)$, we assume the coefficients satisfy $\alpha_{i}\beta_{j,i}\geq 0$ for $p\leq i\leq q$.
This assumption is easily verified in the well-specified setting $\sigma_{*}=\sigma$ [BAGJ21] since $\alpha_{i}=\beta_{i}$, and under link misspecification, it has been directly assumed in prior work [MHWSE23]. We follow [OSSW24] and show that by randomizing the Hermite coefficients of the activation function, a subset of neurons satisfy the above assumption for any degree-$q$ polynomial link function $\sigma_{*}$.
@@
The proof is found in Appendix B.1. Note that in our construction of activation functions for both assumptions, we do not exploit knowledge of the link function $\sigma_{*}$ other than its degree $q$ which decides the constant $C_{\sigma}$. The next theorem shows that by running Phase I for $\tilde{\Theta}(d)$ more steps, a subset of neurons can achieve almost perfect overlap with the index features.
@@
For student neurons with activation satisfying Assumptions 2 and 3 and parameter $\boldsymbol{w}_{j}$ starting from a nontrivial overlap $c$ specified in Theorem 1, for suitable hyperparameters $\eta^{t}=\tilde{O}_{d}(Nd^{-1})$ and $\xi_{j}^{t}=1-\tilde{O}(\varepsilon)$, if we continue to run Phase I of Algorithm 1 for $2T_{1,2}=\tilde{\Theta}_{d}(d\varepsilon^{-2})$ steps, then we achieve $\big{\langle}\boldsymbol{w}_{j}^{2(T_{1,1}+T_{1,2})},\boldsymbol{\theta}\big{\rangle}>1-\varepsilon$ with high probability.
Combining Theorem 1,2 and Lemma 3, we know that after $T_{1}=2(T_{1,1}+T_{1,2})$ steps, there exists some constant $C(q)$ depending only on the degree $q$ such that $1/(C(q)\cdot\mathrm{polylog}(d))$ fraction of the neurons become $\varepsilon$-close to the target direction $\boldsymbol{\theta}$. The following proposition shows that after strong recovery, training the second-layer parameters in Phase II is sufficient for the NN model (3.1) to achieve small generalization error.
@@
$\mathbb{E}_{\boldsymbol{x}}[(f_{\hat{\boldsymbol{\Theta}}}(\boldsymbol{x})-f_{*}(\boldsymbol{x}))^{2}]\lesssim\varepsilon.$
with probability 1 as $d\to\infty$, if we set $T_{2}=C(q)N^{4}\mathrm{polylog}(d)\varepsilon^{-2}$, $N=C(q)\mathrm{polylog}(d)\varepsilon^{-1}$ for some constant $C(q)$ depending on the target degree $q$.
@@
$n=T_{1}+T_{2}\asymp C(q)\cdot(d\varepsilon^{-2}\vee\varepsilon^{-6})\mathrm{polylog}(d),\quad N\asymp C(q)\cdot\varepsilon^{-1}\mathrm{polylog}(d),$ (3.2)
where constant $C(q)$ only depends on the target degree $q$ (although exponentially). Hence we may set $\varepsilon^{-1}\asymp\mathrm{polylog}d$ to conclude an almost-linear sample and computational complexity for learning arbitrary single-index polynomials up to $o_{d}(1)$ population error. This establishes the informal theorem in Section 1.
@@ 4.1 Monomial Transformation Reduces Information Exponent
To prove the main theorem, we first establish the existence of nonlinear label transformation that $(i)$ reduces the information exponent, and $(ii)$ can be easily extracted from SGD updates. If we ignore desideratum $(ii)$, then for polynomial link functions, transformations that decrease the information exponent to at most $2$ have been constructed in [CM20, Section 2.1]. However, prior results are based on the thresholding function with specific offset, and it is not clear if such function naturally arises from SGD with batch reuse. The following proposition shows that the effect of thresholding can also be achieved by a simple monomial transformation.
@@
Let $f:\mathbb{R}\to\mathbb{R}$ be any polynomial with degree up to $p$ and $\mathopen{}\mathclose{{}\left\|{f}}\right\|_{L^{2}(\gamma)}^{2}=1$, then
@@
    There exists some $i\leq C_{q}\in\mathbb{N}_{+}$ such that ${\mathrm{IE}}(f^{i})\leq 2$, where constant $C_{q}$ only depends on $q$.
@@
    Let $f^{\mathrm{odd}}:\mathbb{R}\to\mathbb{R}$ be the odd part of $f$ with $\mathbb{E}_{t\sim\mathcal{N}(0,\boldsymbol{I}_{d})}[f^{\mathrm{odd}}(t)^{2}]\geq\rho>0$. Then there exists some $i\leq C_{q,\rho}\in\mathbb{N}_{+}$ such that ${\mathrm{IE}}(f^{i})=1$, where constant $C_{q,\rho}$ only depends on $q$ and $\rho$.
We make the following remarks.
  • The proposition implies that for any polynomial link function that is not even, there exists some power $i\in\mathbb{N}_{+}$ only depending on the degree of $\sigma_{*}$ such that raising the function to the $i$-the power reduces the information exponent to $1$. For even link functions, the information exponent after arbitrary transformation is at least $2$, and this lowest value can also be achieved by monomial transformation. Furthermore, we provide a uniform upper-bound on the required degree of transformation $i$ via a compactness argument.

### v2

@@ 2.3 Challenges in Establishing Statistical Guarantees
Importantly, the analysis in [DTA+24] does not lead to concrete learnability guarantees for the class of single-index polynomials for the following reasons: $(i)$ it is not clear if an appropriate nonlinear transformation that lowers the information exponent can always be extracted from SGD with reused data, and $(ii)$ the weak recovery guarantee may not translate to a sample complexity for the trained NN to achieve small generalization error. We elaborate these technical challenges below.
@@
To show weak recovery, [DTA+24, Definition 3.1] assumed that the student activation $\sigma$ can reduce the information exponent of the labels to $1$; while a few examples are given, the existence of such transformations in SGD is not guaranteed:
  • The label transformation employed in prior SQ algorithms [CM20] is based on thresholding, which reduces the information exponent to $2$ for any polynomial $\sigma_{*}$; however, isolating such function from SGD updates on the squared loss is challenging. Instead, we make use of monomial transformation which can be extracted from SGD via Taylor expansion.
  • If the link function satisfies $p_{*}\geq 2$, its information exponent after arbitrary nonlinear transformation is at least $2$; such functions are predicted not be not learnable by SGD in the $n\asymp d$ regime [DTA+24]. To handle this setting, we analyze the SGD update up to $\mathrm{poly}(d)$ time, at which a nontrivial overlap can be established by a Grönwall-type argument similar to [BAGJ21]. For $p_{*}=2$, this recovers results on phase retrieval when $\sigma_{*}(z)=z^{2}$ which requires $n=\Omega(d\log d)$ samples.
@@
Note that weak recovery (i.e., $|\langle\boldsymbol{w},\boldsymbol{\theta}\rangle|>\varepsilon$ for some small constant $\varepsilon>0$) is generally insufficient to establish low generalization error of the trained NN. Therefore, we need to show that starting from a nontrivial overlap, subsequent gradient steps can achieve strong recovery of the index features (i.e., $|\langle\boldsymbol{w},\boldsymbol{\theta}\rangle|>1-\varepsilon$), despite the link misspecification. After the first-layer parameters align with the target function, we train the second-layer parameters with SGD to learn the link function $\sigma_{*}$ with the aid of random bias units [DLS22].
3 Learning Polynomial $f_{*}$ in Linear Sample Complexity
We first consider the setting where $\sigma_{*}$ is polynomial with degree $q$ specified as follows.
@@ 3 Learning Polynomial $f_{*}$ in Linear Sample Complexity
For single-index polynomials, we do not expect a computational-to-statistical gap under the SQ class [CM20] — indeed, we will establish learning guarantees near the information theoretic limit $n\asymp d$.
@@ 3.1 Training Algorithm
We train the following two-layer network with $N$ neurons using SGD to minimize the squared loss:
@@
  • The most crucial part in Phase I of Algorithm 1 is the reuse of the same minibatch in the gradient computation. Specifically, we sample a fresh batch of training examples in every two GD steps; this enables us to extract non-CSQ terms from two consecutive gradient updates outlined in (2.3).
  • We introduce an interpolation step between the current and previous iterates with hyperparameter $\xi$ to stabilize the training dynamics; this resembles a negative momentum often seen in optimization algorithms [AZ18, ZLBH19]; the role of this interpolation is discussed in Section 4.2. We use a projected gradient update ~$\tilde{\nabla}_{\boldsymbol{w}}\mathcal{L}(\boldsymbol{w})=(\boldsymbol{I}_{d}-\boldsymbol{w}^{2t}{\boldsymbol{w}^{2t}}^{\top})\nabla_{\boldsymbol{w}}\mathcal{L}(\boldsymbol{w})$ for steps $2t$ and $12t+$ 1, where $\nabla_{\boldsymbol{w}}$ is the Euclidean gradient; similar use of projection also appeared in [DNGL23, AAM23].
@@
Input : Step sizes $\eta^{t}$; momentum parameters $\xi^{t}$; training time $T_{1},T_{2}$; $\ell_{2}$ regularization $\lambda$.
@@
3Initialize $\boldsymbol{w}^{0}_{j}\sim\mathbb{S}^{d-1}(1)$, $a_{j}\sim\mathrm{Unif}\{\pm c_{a}\}$.
@@
6             if $t$ is even then
                   $\boldsymbol{x}\sim\mathcal{N}(0,\boldsymbol{I}_{d}),~{}y=f_{*}(\boldsymbol{x})+\varsigma$ ;
                    // Draw i.i.d. data $(\boldsymbol{x},y)$
                   ${\boldsymbol{w}}^{t}_{j}\leftarrow\boldsymbol{w}^{t}_{j}-\xi_{j}^{t}(\boldsymbol{w}^{t}_{j}-\boldsymbol{w}^{t-2}_{j})$,   (when $t>0$) ;
                    // Interpolation step
                   $\boldsymbol{w}^{t}_{j}\leftarrow\boldsymbol{w}^{t}_{j}/\|\boldsymbol{w}^{t}_{j}\|$ ;
                    // Normalization
7
            $\boldsymbol{w}^{t+1}_{j}\leftarrow\boldsymbol{w}^{t}_{j}-\eta^{t}\tilde{\nabla}_{\boldsymbol{w}}(f_{\boldsymbol{\Theta}}(\boldsymbol{x})-y)^{2},~{}~{}(j=1,\dots,N)$ ;
              // SGD step
8
9
10
11Initialize $b_{j}\sim\mathrm{Unif}([-C_{b},C_{b}])$.
12Phase II: SGD on second-layer parameters
       $\hat{\boldsymbol{a}}\leftarrow\mathrm{argmin}_{\boldsymbol{a}\in\mathbb{R}^{N}}\frac{1}{T_{2}}\sum_{i=1}^{T_{2}}\mathopen{}\mathclose{{}\left(f_{\boldsymbol{\Theta}}(\boldsymbol{x}_{i})-y_{i}}\right)^{2}+\lambda\|\boldsymbol{a}\|^{2}$ ;
        // Ridge regression
13
14
@@
3.2 Convergence and Sample Complexity
@@
We require the activation function to be a polynomial $\sigma(z)=\sum_{i=0}^{C_{\sigma}}\beta_{i}\mathsf{He}_{i}(z)$ and its degree $C_{\sigma}$ to be sufficiently large so that $C_{\sigma}\geq C_{q}$ holds ($C_{q}$ is defined in Proposition 6). For all $2\leq\ell\leq C_{\sigma}$ and $k=0,1$, we assume that $H\big{(}\sigma^{(\ell)}(\sigma^{(1)})^{\ell-1};k\big{)}>0$.
As discussed in Appendix B.1, for a given $\sigma_{*}$, the above assumption only needs to be met for one pair of $(k,\ell)$. Appendix B.1.3 states that $H\big{(}\sigma^{(\ell)}(\sigma^{(1)})^{\ell-1};k\big{)}\neq 0$ also suffices if we set the momentum parameter $\xi$ differently. Now we verify this condition for a wide range of polynomial activations.
@@
Given $\ell\geq 2$ and $k\geq 0$. For $C_{\sigma}\geq\frac{2\ell+k-1}{\ell}$, if we choose $\{\beta_{i}\}_{i=0}^{C_{\sigma}}$ where $\beta_{i}$ is randomly drawn from some non-empty interval $[a_{i},b_{i}]$, then $H(\sigma^{(\ell)}(\sigma^{(1)})^{\ell-1};k)\neq 0$ with probability 1.
The next theorem states that $n=\tilde{\Theta}(d)$ samples are sufficient for SGD to achieve weak recovery.
@@
Under Assumptions 1 and 2, for suitable choices of hyperparameters $\eta^{t}=\tilde{O}_{d}(Nd^{-1})$ and $1-\xi^{t}=o_{d}(1)$, there exists constant $C(q)$ such that after Phase I of Algorithm 1 is run for $2T_{1,1}=C(q)\cdot d\mathrm{polylog}(d)$ steps, with high probability, there exists a subset of neurons $\boldsymbol{w}_{j}^{2T_{1}}\in\mathcal{W}$ with $|\mathcal{W}|=\tilde{\Theta}(N)$ such that $\big{|}\langle\boldsymbol{w}_{j}^{2T_{1}},\boldsymbol{\theta}\rangle\big{|}>c$ for some $c\gtrsim 1/\mathrm{polylog}(d)$.
@@
Given the Hermite expansions $\sigma_{*}(z)=\sum_{i=p}^{q}\alpha_{i}\mathsf{He}_{i}(z)$, $\sigma_{j}(z)=\sum_{i=0}^{C_{\sigma}}\beta_{j,i}\mathsf{He}_{i}(z)$, we assume the coefficients satisfy $\alpha_{i}\beta_{j,i}\geq 0$ for $p\leq i\leq q$.
This assumption is easily verified in the well-specified setting $\sigma_{*}=\sigma$ [BAGJ21] since $\alpha_{i}=\beta_{i}$, and under link misspecification, it has been directly assumed in prior work [MHWSE23]. We follow [OSSW24a] and show that by randomizing the Hermite coefficients of the activation function, a subset of neurons satisfy the above assumption for any degree-$q$ polynomial link function $\sigma_{*}$.
@@
Note that in our construction of activation functions for both assumptions, we do not exploit knowledge of the link function $\sigma_{*}$ other than its degree $q$ which decides the constant $C_{\sigma}$; see Appendix B.1 for more discussion of Assumption 3 and Lemma 4. The next theorem shows that by running Phase I for $\tilde{\Theta}(d)$ more steps, a subset of neurons achieves sufficiently large overlap with the index features.
@@
For student neurons satisfying Assumptions 2, 3 and parameter $\boldsymbol{w}_{j}$ starting from nontrivial overlap $c>0$ specified in Theorem 1, if Phase I of Algorithm 1 continues for $2T_{1,2}=\tilde{\Theta}_{d}(d\varepsilon^{-2})$ steps with hyperparameters $\eta^{t}=\tilde{O}_{d}(Nd^{-1}\varepsilon)$, $\xi^{t}=1$, we achieve $\big{\langle}\boldsymbol{w}_{j}^{2(T_{1,1}+T_{1,2})},\boldsymbol{\theta}\big{\rangle}>1-\varepsilon$ with high probability.
The following proposition shows that after strong recovery, training the second-layer parameters in Phase II is sufficient for the NN model (3.1) to achieve small generalization error.
@@
$\mathbb{E}_{\boldsymbol{x}}[(f_{\hat{\boldsymbol{\Theta}}}(\boldsymbol{x})-f_{*}(\boldsymbol{x}))^{2}]\lesssim\varepsilon^{2}.$
with probability 1 as $d\to\infty$, if we set $T_{2}=C(q)N^{4}\mathrm{polylog}(d)\varepsilon^{-4}$, $N=C(q)\mathrm{polylog}(d)\varepsilon^{-1}$ for some constant $C(q)$ depending on the target degree $q$.
@@
$n=T_{1}+T_{2}\asymp C(q)\cdot(d\varepsilon^{-2}\vee\varepsilon^{-8})\mathrm{polylog}(d),\quad N\asymp C(q)\cdot\varepsilon^{-1}\mathrm{polylog}(d),$ (3.2)
where constant $C(q)$ only depends on the target degree $q$ (although exponentially). Hence we may set $\varepsilon^{-1}\asymp\mathrm{polylog}d$ to conclude an almost-linear sample and computational complexity for learning arbitrary single-index polynomials up to $o_{d}(1)$ population error.
@@ 4.1 Monomial Transformation Reduces Information Exponent
To prove the main theorem, we first establish the existence of nonlinear label transformation that $(i)$ reduces the information exponent, and $(ii)$ can be easily extracted from SGD updates. If we ignore desideratum $(ii)$, then for polynomial link functions, transformations that decrease the information exponent to at most $2$ have been constructed in [CM20, Section 2.1]. However, prior results are based on the thresholding function, and it is not clear if such function naturally arises from SGD with batch reuse. The following proposition shows that the effect of thresholding can also be achieved by a simple monomial transformation where the required degree can be uniformly upper bounded.
@@
Let $g:\mathbb{R}\to\mathbb{R}$ be any polynomial with degree up to $p$ and $\mathopen{}\mathclose{{}\left\|{g}}\right\|_{L^{2}(\gamma)}^{2}=1$, then
@@
    There exists some $i\leq C_{q}\in\mathbb{N}_{+}$ such that ${\mathrm{IE}}(g^{i})\leq 2$, where constant $C_{q}$ only depends on $q$.
@@
    Let $g^{\mathrm{odd}}:\mathbb{R}\to\mathbb{R}$ be the odd part of $g$ with $\|g^{\mathrm{odd}}\|_{L^{2}(\gamma)}^{2}\geq\rho>0$. Then there exists some $i\leq C_{q,\rho}\in\mathbb{N}_{+}$ such that ${\mathrm{IE}}(g^{i})=1$, where constant $C_{q,\rho}$ only depends on $q$ and $\rho$.
The proof can be found in Appendix A. We make the following remarks.
  • The proposition implies that for any polynomial link function that is not even, there exists $i\in\mathbb{N}_{+}$ only depending on the degree of $\sigma_{*}$ such that raising the function to the $i$-th power reduces the information exponent to $1$ ($p_{*}=1$). For even $\sigma_{*}$, the information exponent after arbitrary transformation is at least $2$ ($p_{*}=2$), which can also be attained by monomial transformation. Furthermore, we provide a uniform upper-bound on the required degree of transformation $i$ via a compactness argument.